# Colloid-like scale-free cluster-cluster aggregation during polymer collapse

### v1

@@ Abstract
An extended polymer collapses to form a globule when subjected to a quench below the collapse transition temperature. The process begins with the formation of clusters of monomers or “pearls”. The nascent clusters merge, resulting in growth of the average cluster size $C_{s}$, eventually leading to a single globule. The aggregation of the clusters are known to be analogous to droplet coalescence. This suggests a striking resemblance between such an aggregation and cluster-cluster aggregation in colloidal self-assembly, which is characterized by a universal dynamic scaling behavior. Motivated by that, here, we verify the presence of such dynamic scaling during the collapse of a polymer with varying bending stiffness $\kappa$, using molecular dynamic simulations. We probe the dynamics via time evolution of the size distribution of clusters $N_{s}(t)$ and growth of $C_{s}(t)$. Irrespective of $\kappa$, we observe the power-law scalings $C_{s}(t)\sim t^{z}$ and $N_{s}(t)\sim t^{-w}s^{-\tau}$, of which only the cluster growth is universal with $z\approx 1.65$. Importantly, our results indeed show that $N_{s}(t)$ exhibits a dynamic scaling of the form $N_{s}(t)\sim s^{-2}f(s/t^{z})$, indicative of a scale-free cluster growth. Interestingly, for flexible and weakly stiff polymers the dynamic exponents obey the relation $w=2z$, as also found in diffusion-controlled cluster-cluster aggregation of particles. For $\kappa\geq 5$, the exponents show deviation from this relation, which grows continuously with $\kappa$. We identify the differences in local structures of the clusters formed, leading to variations in cluster-size dependence of the effective diffusion constant to be the origin of the above deviation. We also discuss potential experimental strategies to directly visualize the observed dynamic scaling in a collapsing polymer.
@@ I Introduction
On the other hand, as already mentioned there have been multiple studies that indicate the presence of cluster-cluster aggregation during the collapse of a polymer. In spite of that, to the best of our knowledge, no effort has been dedicated to apply the Vicsek-Family scaling principles in understanding the same. The necessity of such an approach, advocating universality in cluster sizes is increasingly becoming important as conformational changes in complex bio-macromolecule like proteins require better understanding of the underlying principles of polymer physics. To this end, here, by means molecular dynamics (MD) simulations we investigate the kinetics of collapse of homopolymers, ranging from flexible to moderately stiff ones (semiflexible), with an emphasis on exploring the presence of any dynamic scaling, as found in colloidal systems.
@@ II Model and Methods
The modified potential behaves qualitative the same as $V_{\text{LJ}}$, however, is continuous and differentiable at $r=r_{c}$. The energy penalty for bending of the polymer backbone is given by
@@
Since we study the nonequilibrium dynamics of the collapse transition of a polymer, an extended-coil conformation of a polymer is desired as the initial configuration. For that we use a self-avoiding walk of length $N-1$ as a polymer with $N$ monomers. Using this self-avoiding walk we perform MD simulation at a high $T$ for an extended period such that polymer chain attains an extended-coil conformation. Following that we start the nonequilibrium evolution by setting $T=1$. We present results for polymers with three different chain lengths, $N=512$, $1024$ and $2048$. As we are interested in the kinetics of collapse transition, here, we present result for polymer with moderate bending stiffness, i.e., $\kappa\in[0,10]$. For even larger $\kappa$, the collapse transition accompanies with other complex conformational transitions Majumder et al. (2021). Except the evolution snapshots and contact maps all data presented are averaged over $300$ independent realizations obtained by using different seed for the random number while generating the self-avoiding walk.
@@ IV.1 Flexible polymers
We begin by showing representative conformations of a flexible polymer illustrating the time evolution during its collapse, in Fig. 1. Starting from an extended-coil-like conformation, at early times, clusters of monomers or “pearls” develop randomly along the chain. Such developments result in appearance of locally dense regions. To identify these clustered regions on the chain, we estimate the inter-monomer pairwise distances for non-bonded monomers. We consider a pair of monomers to be in contact, if their distance is less than $2.5\sigma$. On the two-dimensional contact maps, the colored diagonal corresponds to self contacts. These are almost the only contacts present at early times as shown in the first frame. Transient proximity of different segments of the fluctuating polymer generates the sporadic off-diagonal contacts. The developments of clusters with time appear as few off-diagonal contacts. The second and the third frame depict such intermediate stage of collapse representing the “pearl-necklace”. Visual inspection of the snapshot as the collapse progresses suggests decreased number of clusters, but with increased sizes . Consistent with the prediction of the pearl-necklace representation of the collapse, inspection of the trajectories reveal that the nucleated clusters merge with each other forming larger clusters. The corresponding contact map demonstrates this growth by increase in size of regions of off-diagonal or long-range contacts at the expense of their number. Finally, a single globule is formed after aggregation of all the clusters (final frame in Fig. 1), resulting in appearance of contacts throughout the chain.
[…]
@@
Figure 3: Cluster growth in flexible polymers. (a) Growth of the average cluster size $\langle C_{s}(t)\rangle$ during the collapse of polymers of three different lengths $N$. The dashed line shows the consistency of the data with a power-law growth with an exponent $z=1.65$. (b) Plots of the scaling functions $Y(y)$, demonstrating the collapse of data for different $N$ through a finite-size scaling analysis providing an unambiguous estimate of $z$. In both (a) and (b) the data are shown on a double-log scale.
@@
After identifying the time regime of interest, we explore whether the dynamic scaling in Eq. (8) or (11) can be extended to aggregations of clusters of monomers under the topological constraint of a polymer chain. As the first step, we check the variation of the average cluster size with time following Eq. (6) [see Fig. 3(a)]. The rise of $\langle C_{s}(t)\rangle$ for different $N$ follow similar profiles, barring the saturation at late times. Such deviations are due to effects of finite lengths of the polymer chains. The preceding similarity suggests presence of identical kinetics of growth. The sluggish growth of $\langle n_{c}(t)\rangle$ for small $t$ ($\lessapprox 10$) is a signature of monomer acquisition by the existing clusters or “pearls”. This stage is followed by a faster growth which exhibits a linear behavior on the double-log scale, consistent with a power-law behavior in Eq. (7) with $z=1.65$. A comparison of time scales between Figs. 2 and 3(a) confirms that the time period when $\langle C_{s}(t)\rangle$ shows increase is indeed the cluster-cluster aggregation regime, i.e., our regime of interest.
@@
$y=\frac{\langle C_{s}\rangle-C_{0}}{N-C_{0}},$ (18)
in this exercise by tuning the value $z$, one thus expects that the data for different $N$ to collapse onto each other for the appropriate choice of $z$ along with a $Y(y)\sim y^{-z}$ behavior in the finite-size unaffected regime. The optimum result of this exercise is demonstrated in Fig. 3(b) that confirms the value of $z\approx 1.65$.
Figure 4: Kinetics of clusters of fixed sizes in flexible polymers. Time dependence of the number $N_{s}$ of clusters of fixed size $s$ on a double-log scale, during the collapse. Plots for multiple $s$ across different $N$ are included. The abscissa is scaled with $N^{1/z}$ for a better visualization. $N_{s}(t)$ display both monotonic and bell-shaped profiles.
@@
Next, we measure the variation in the number of clusters $N_{s}(t)$ of a fixed size $s$ with time, shown in Fig. 4. We choose the values of $s\geq 20$, which belong to the scaling regime in $\langle C_{s}(t)\rangle$ [see Fig. 3(a)]. The plots include results from three different $N$ to distinctly identify the finite-size unaffected regime. The $t$ axis is scaled with $N^{1/z}$ to facilitate a better visualization of data for different $N$. At early times, mostly small clusters are present. Whereas, clusters with large $s$ appear later at the expense of the smaller ones. Thus, $N_{s}(t)$ for smaller $s$ starts the decay from a higher value. Also, their decay are concurrent with the rise in the number of larger clusters. In the intermediate to late times, data for different $s$ across different $N$ collapse on a power-law decay function. Keeping in mind Eq. (9), in this regime we fit the expression
$N_{s}(t)=A_{w}t^{-w},$ (19)
which yields $w\approx 3.02\pm 0.12$.
Having both $z$ and $w$ in hand, we move on to investigate the distribution of cluster sizes for different times in the scaling regime and check if there exists any Viscek-Family-like dynamic scaling. For that in Fig. 5 we present the scaling function $f(x)\equiv s^{2}N_{s}(t)$ as a function of the scaling variable $x\equiv s/t^{z}$ for different $t$ using data for a polymer with $N=2048$. The data for different $t$ show a reasonable data collapse. The data in the regime $x\ll 1$ show an increase consistent with the power law $\sim x^{2}$ as shown by the solid line, which indicates that possibly $\tau=0$. Thus one can expect that the validity of the scaling relation Eq. (12). Recalling $z=1.65(3)$ and $w=3.12(12)$, it can be inferred that within numerical error for flexible polymer indeed that is the case. Overall, we conclude that colloid-like dynamic scaling for cluster-cluster aggregation can also be realized in case of clusters composed of monomers during collapse of a flexible polymer.
@@ IV.2 Semi-flexible polymers
Representative evolution snapshots and corresponding contact maps during the collapse of a semiflexible polymer with a reasonably high value of bending stiffness $\kappa=8$ are presented in Fig. 6. At early times, there is development of clusters, leading to few contacts close to the diagonal. This is followed by aggregation of the small clusters forming of larger regions of off-diagonal contacts. Eventually, a globule is formed where long-range contacts are present over the whole chain. The progressions are similar to those in flexible polymers. However, significantly more order with diamond-like moieties can be noticed in the contact maps for the semi-flexible case. Such order has been reported earlier, and the diamond-like structures in contact maps have been attributed to the ordered local structures for semi-flexible polymers Lappala and Terentjev (2013).
@@
Having an idea about the phenomenology of cluster-cluster aggregation during collapse of semiflexible polymers, we move on to quantify the cluster growth as a function of time. Corresponding plots for different $\kappa$ for a fixed chain length $N=2048$ are presented in Fig. 7(a). As in the case of flexible polymers, $\langle C_{s}(t)\rangle$ for all $\kappa$ display a relatively slow rise at early time, followed by a power-law growth in the scaling regime. As $\kappa$ increases the entry into the scaling regime gets delayed. This can be further confirmed from corresponding plots for time dependence of the number of clusters $n_{c}(t)$ for $\kappa=5$ and $\kappa=8$, presented in the inset of Fig. 7(a). There, the non-monotonic profile for $\kappa=8$ appears because of the time-independent size cut-off, $s_{c}=10$ used to identify the clusters. If one compares the data for the cluster growth for different $\kappa$, it appears that in the scaling regime they are almost parallel to each other, implying the presence of a universal growth exponent $z$, irrespective of $\kappa$. Nevertheless, we fit the ansatz in Eq. (16) to the data, results of which are compiled in Table 1. The results yield no systematic dependence of the estimated $z$ values on $\kappa$. Rather it fluctuates around a mean $z=1.65$. Of course, to further confirm the presence of a $\kappa$-independent $z$, one can rely on finite-size scaling analysis of individual $\kappa$, as done for the flexible case in the previous subsection. However, for the sake of brevity, here, we do not present such exercises for different $\kappa$.
Table 1: Fitting of the ansatz in Eq. (16) to the data for time dependence of the cluster size for polymers with different stiffness $\kappa$.
$\kappa$ $C_{0}$ $A_{z}$ $z$ $t$-range
00 $32.54(10)$ $0.37(1)$ $1.65(3)$ $[1:140]$
$1$ $18.65(44)$ $0.22(1)$ $1.66(2)$ $[1:140]$
$2$ $15.05(10)$ $0.20(2)$ $1.61(2)$ $[1:150]$
$3$ $12.93(52)$ $0.16(2)$ $1.60(2)$ $[1:200]$
$4$ $8.56(10)$ $0.12(2)$ $1.63(3)$ $[20:300]$
$5$ $6.97(50)$ $0.10(1)$ $1.66(3)$ $[20:300]$
$6$ $4.80(20)$ $0.08(8)$ $1.67(1)$ $[30:320]$
$7$ $3.07(20)$ $0.07(1)$ $1.68(2)$ $[30:340]$
$8$ $3.27(10)$ $0.05(1)$ $1.69(3)$ $[30:340]$
$9$ $3.20(20)$ $0.05(2)$ $1.63(2)$ $[30:360]$
$10$ $3.10(20)$ $0.03(3)$ $1.61(3)$ $[30:360]$
@@
Figure 8: Verification of dynamic scaling in semiflexible polymers. Double-log plot of the scaling function $f(x)\equiv\langle s^{2}N_{s}(t)\rangle$ against the scaling variable $x\equiv s/t^{z}$ at different times for a polymer of length $N=2048$ with bending stiffness (a) $\kappa=5$ and (b) $\kappa=8$. The dashed black lines show the consistency of the data with the power law $\sim x^{\delta}$ in the regime $x\ll 1$, with respective values of $\delta$ as indicated. Collapse of data at different times confirm the presence of colloid-like dynamic scaling in semiflexible polymers.
@@
Table 2: Fitting of the ansatz in Eq. (19) to the data for decay of the number of cluster of a fixed size $s$ as a function of time for polymers with fixed $N=2048$ and different $\kappa$.
$\kappa$ $A_{w}(10^{4})$ $w$ range
00 $8.14(37)$ $3.02(12)$ $[30,150]$
$1$ $4.39(93)$ $3.21(6)$ $[30,150]$
$2$ $2.70(77)$ $2.98(7)$ $[50,250]$
$3$ $2.07(57)$ $2.70(6)$ $[30,100]$
$4$ $38.68(57)$ $2.87(9)$ $[50,200]$
$5$ $52.35(75)$ $2.88(5)$ $[50,200]$
$6$ $32.92(90)$ $2.70(10)$ $[50,200]$
$7$ $7.44(85)$ $2.28(8)$ $[50,200]$
$8$ $4.96(10)$ $2.08(9)$ $[80,300]$
$9$ $7.93(76)$ $2.07(9)$ $[80,300]$
$10$ $3.32(98)$ $1.82(10)$ $[100,400]$
@@
To obtain the exponent $w$, we probe the time dependence of $N_{s}(t)$, i.e., by calculating the variations in the count of clusters of size $s$ with time. Representative plots for $N=2048$ and $s=40$ is presented in Fig. 7(b). As $\kappa$ increases one notices a progressively prominent bell-shaped non-monotonic behavior for the same value of $s$. This is due to the fact that with increasing $\kappa$ nucleation of clusters to form stable clusters gets delayed, as also evident from pots of number of clusters $\langle n_{c}(t)\rangle$ in the inset of Fig. 7(a). For each considered $\kappa$, we fit the region with a common decay for different $s$ (not shown here) to the expression in Eq. (19). The obtained estimates of $w$ are quoted in Table 2. The value of $w$ appears to decrease with increasing $\kappa$. Also, the value of $w$ becomes smaller than $2z$ as the polymer becomes stiffer. This suggests deviation from the scaling relation in of Eq. (12). The deviations motivate us to ask whether the dynamic scaling at all exists for semi-flexible polymers. If yes, then possibly the cluster-cluster aggregation in semiflexible polymer can be represented by the other kind of dynamic scaling functional form as embedded in Eqs. (8) and (10).
@@
Table 3: Fitting of Eq. (20) to the data of the scaling function $f(x)$ as a function of the scaling variable $x$ at different times for polymers with fixed length $N=2048$ and different $\kappa$.
$\kappa$ $A_{\delta}(10^{5})$ $\delta$
00 $2.18(37)$ $2.20(25)$
$1$ $4.50(17)$ $2.12(18)$
$2$ $3.30(77)$ $2.21(21)$
$3$ $2.53(57)$ $2.16(17)$
$4$ $3.07(33)$ $2.11(13)$
$5$ $17.33(75)$ $1.92(20)$
$6$ $16.56(90)$ $1.83(10)$
$7$ $13.01(85)$ $1.71(8)$
$8$ $13.26(10)$ $1.68(10)$
$9$ $5.20(26)$ $1.52(10)$
$10$ $5.96(28)$ $1.54(12)$

### v2

# Scale-free cluster-cluster aggregation during polymer collapse

@@ Abstract
Abstract: An extended polymer collapses to form a globule when subjected to a quench below the collapse transition temperature. The process begins with the formation of clusters of monomers or “pearls”. The nascent clusters merge, resulting in growth of the average cluster size $C_{s}$, eventually leading to a single globule. The aggregation of the clusters are known to be analogous to droplet coalescence. This suggests a striking resemblance between such an aggregation and cluster-cluster aggregation found in many particle systems, like in colloidal self-assembly, typically characterized by a universal dynamic scaling behavior. Motivated by that, here, we verify the presence of such dynamic scaling during the collapse of a polymer with varying bending stiffness $\kappa$, using molecular dynamic simulations. We probe the dynamics via time evolution of the size distribution of clusters $N_{s}(t)$ and growth of $C_{s}(t)$. Irrespective of $\kappa$, we observe the power-law scalings $C_{s}(t)\sim t^{z}$ and $N_{s}(t)\sim t^{-w}s^{-\tau}$, of which only the cluster growth is universal with $z\approx 1.67$. Importantly, our results indeed show that $N_{s}(t)$ exhibits a dynamic scaling of the form $N_{s}(t)\sim s^{-2}f(s/t^{z})$, indicative of a scale-free cluster growth. Interestingly, for flexible and weakly stiff polymers the dynamic exponents obey the relation $w=2z$, as also found in diffusion-controlled cluster-cluster aggregation of particles. For $\kappa\geq 5$, the exponents show deviation from this relation, which grows continuously with $\kappa$. We identify the differences in local structures of the clusters formed, leading to variations in cluster-size dependence of the effective diffusion constant to be the origin of the above deviation. We also discuss potential experimental strategies to directly visualize the observed dynamic scaling in a collapsing polymer.
@@ I Introduction
On the other hand, as already mentioned there have been multiple studies that indicate the presence of cluster-cluster aggregation during the collapse of a polymer. In spite of that, to the best of our knowledge, no effort has been dedicated to apply the Vicsek-Family scaling principles in understanding the same. The necessity of such an approach, advocating universality in cluster sizes is increasingly becoming important as conformational changes in complex bio-macromolecule like proteins require better understanding of the underlying principles of polymer physics. To this end, here, by means of molecular dynamics (MD) simulations we investigate the kinetics of collapse of homopolymers, ranging from flexible to moderately stiff ones (semiflexible), with an emphasis on exploring the presence of any dynamic scaling, as found in colloidal systems.
@@ II Model and Methods
The modified potential behaves qualitatively the same as $V_{\text{LJ}}$, however, is continuous and differentiable at $r=r_{c}$. The energy penalty for bending of the polymer backbone is given by
@@
Since we study the nonequilibrium dynamics of the collapse transition of a polymer, an extended-coil conformation of a polymer is desired as the initial configuration. For that we use a self-avoiding walk of length N1N-1 as a polymer with $N$ monomers. Using this self-avoiding walk we first perform MD simulation at a high $T$ for a period $\tau_{\rm eq}\sim N^{2}$ such that the polymer chain attains an extended-coil conformation. Following that we start the nonequilibrium evolution by setting $T=1$. As we are interested in the kinetics of collapse transition, here, we present result for polymer with moderate bending stiffness, i.e., $\kappa\in[0,10]$. For even larger $\kappa$, the collapse transition accompanies with other complex conformational transitions Majumder et al. (2021). Majority of the results are based on polymers of chain lengths, $N\leq 2048$. However, for further confirmation, for three specific values of $\kappa$, we present results from relatively much longer chain length of $N=8192$. The equilibration time $\tau_{\rm eq}$ at high $T$ for $N=512$ is $100\tau_{0}$. For other chain lengths, $\tau_{\rm eq}$ is chosen using the relation $\tau_{\rm eq}\sim N^{2}$. Following the quench, we simulated the chain until it collapses to a single globule. The typical timescale for the collapse is $\tau_{c}\sim N$ in MD simulations Majumder et al. (2020). In this study, as we will see later that the collapse time is also depended on the polymer stiffness. Typical maximum run time chosen for a chain of length $N=2048$ varies from $200\tau_{0}$ to $800\tau_{0}$ as $\kappa$ varies from 0 to $10$. Except the evolution snapshots and contact maps all data presented are averaged over $500$ independent realizations obtained by using different seed for the random number while generating the self-avoiding walk.
@@ III Dynamic scaling theory
Origin of the dynamic scaling of the form (8) or (11) was purely empirical Vicsek and Family (1984); Meakin and Vicsek (1985). The scaling relation in Eq. (14) is purely a consequence of the constant density of the system. Thus, even though it was initially realized in diffusion-limited aggregation of clusters such as during colloidal self-assembly, over the years it has been observed that systems with different transport mechanisms too exhibit such dynamic scaling. This includes interface growth during roughening transition, active Eden process, growth due to stochastic ballistic deposition and even interface growth in turbulent liquid crystals Family and Vicsek (1985); Plischke and Rácz (1985); Kim and Kosterlitz (1989); Maunuksela et al. (1997); Myllys et al. (2001); Takeuchi and Sano (2010).
@@ IV.1 Flexible polymers
We begin by showing representative conformations of a flexible polymer of length $N=2048$ illustrating the time evolution during its collapse, in Fig. 1. Starting from an extended-coil-like conformation, at early times, clusters of monomers or “pearls” develop randomly along the chain. Such developments result in appearance of locally dense regions. To identify these clustered regions on the chain, we estimate the inter-monomer pairwise distances for non-bonded monomers. We consider a pair of monomers to be in contact, if their distance is less than $2.5\sigma$. On the two-dimensional contact maps, the colored diagonal corresponds to self contacts. These are almost the only contacts present at early times as shown in the first frame. Transient proximity of different segments of the fluctuating polymer generates the sporadic off-diagonal contacts. The developments of clusters with time appear as few off-diagonal contacts. The second and the third frame depict such intermediate stage of collapse representing the “pearl-necklace”. Visual inspection of the snapshot as the collapse progresses suggests decreased number of clusters, but with increased sizes . Consistent with the prediction of the pearl-necklace representation of the collapse, inspection of the trajectories reveal that the nucleated clusters merge with each other forming larger clusters. The corresponding contact map demonstrates this growth by increase in size of regions of off-diagonal or long-range contacts at the expense of their number. Finally, a single globule is formed after aggregation of all the clusters (final frame in Fig. 1), resulting in appearance of contacts throughout the chain.
To quantify these stages of collapse, we present the time evolution of the average number of clusters $\langle n_{c}\rangle$ for three different chain lengths in Fig. 2. Here and later $\langle\dots\rangle$ indicates averaging over different simulation runs performed using independent initial conformations. Following the temperature quench, clusters nucleate fast, resulting in an almost instantaneous rise of $\langle n_{c}\rangle$. Afterwards, for a short time period spanning up to about $t=10$, the profiles remain almost flat. The plots are shifted upwards for longer polymers. A simultaneous inspection of the snapshots and contact maps in Figure 1 suggests that the clusters continue to grow during this time window. Clearly, at this stage the growth occurs by acquiring monomers from the chains connecting the clusters, having no impact on the number of clusters. The time span up to this stage appears to be independent of $N$. A monotonic decay to unity or a collapsed globule state follows the plateau. Evidently, aggregation of smaller clusters result in such decrease of $\langle n_{c}\rangle$. As expected, the formation of globule takes longer for larger $N$. Here, we focus on the spatio-temporal scales, where $\langle n_{c}\rangle$ undergoes the monotonic decay at intermediate to late stages of the collapse, and identify this cluster-cluster aggregation dominated stage as our regime of interest.
@@
Figure 3: Cluster growth in flexible polymers. (a) Growth of the average cluster size $\langle C_{s}(t)\rangle$ during the collapse of polymers of three different lengths NN. The dashed line shows the consistency of the data with a power-law growth with an exponent $z=1.67$. (b) Plots of the scaling functions $Y(y)$, demonstrating the collapse of data for different NN through a finite-size scaling analysis providing an unambiguous estimate of $z$. In both (a) and (b) the data are shown on a double-log scale.
@@
After identifying the time regime of interest, we explore whether the dynamic scaling in Eq. (8) or (11) can be extended to aggregations of clusters of monomers under the topological constraint of a polymer chain. As the first step, we check the variation of the average cluster size with time following Eq. (6) [see Fig. 3(a)]. The rise of $\langle C_{s}(t)\rangle$ for different $N$ follow similar profiles, barring the saturation at late times. Such deviations are due to effects of finite lengths of the polymer chains. The preceding similarity suggests presence of identical kinetics of growth. The sluggish growth of $\langle n_{c}(t)\rangle$ for small $t$ ($\lessapprox 10$) is a signature of monomer acquisition by the existing clusters or “pearls”. This stage is followed by a faster growth which exhibits a linear behavior on the double-log scale, consistent with a power-law behavior in Eq. (7) with $z=1.67$. A comparison of time scales between Figs. 2 and 3(a) confirms that the time period when $\langle C_{s}(t)\rangle$ shows increase is indeed the cluster-cluster aggregation regime, i.e., our regime of interest.
@@
$y=\frac{(N-C_{0})^{1/z}}{t},$ (18)
in this exercise by tuning the value $z$, one thus expects that the data for different $N$ to collapse onto each other for the appropriate choice of $z$ along with a $Y(y)\sim y^{-z}$ behavior in the finite-size unaffected regime. The optimum result of this exercise is demonstrated in Fig. 3(b) that confirms the value of $z\approx 1.67$.
Figure 4: Kinetics of clusters of fixed sizes in flexible polymers. Time dependence of the number $N_{s}$ of clusters of fixed size $s$ on a double-log scale, during the collapse. Plots for multiple $s$ across different NN are included. The abscissa is scaled with $N^{1/z}$ for a better visualization. $N_{s}(t)$ display both monotonic and bell-shaped profiles. The dashed line shows the consistency of the data with the power-law decay.
@@
Next, we measure the variation in the number of clusters $N_{s}(t)$ of a fixed size $s$ with time, shown in Fig. 4. We choose the values of $s\geq 20$, which belong to the scaling regime in $\langle C_{s}(t)\rangle$ [see Fig. 3(a)]. The plots include results from three different $N$ to distinctly identify the finite-size unaffected regime. The $t$ axis is scaled with $N^{1/z}$ to facilitate a better visualization of data for different $N$. At early times, mostly small clusters are present. Whereas, clusters with large $s$ appear later at the expense of the smaller ones. Thus, $N_{s}(t)$ for smaller $s$ starts the decay from a higher value. Also, their decay are concurrent with the rise in the number of larger clusters. In the intermediate to late times, data for different $s$ across different $N$ collapse on a power-law decay function. Keeping in mind Eq. (9), in this regime one can quantify the data using the expression
$N_{s}(t)=A_{w}t^{-w}.$ (19)
The dashed line Fig. 4 show the consistency of the data with this power law decay with $w=3.21$. In the next subsection, we present how the mean exponent $\langle w\rangle$ can be extracted from the data for both flexible and semiflexible polymers.
Having both $z$ and $w$ in hand, we move on to investigate the distribution of cluster sizes for different times in the scaling regime and check if there exists any Vicsek-Family-like dynamic scaling. For that in Fig. 5 we present the scaling function $f(x)\equiv s^{2}N_{s}(t)$ as a function of the scaling variable $x\equiv s/t^{z}$ for different $t$ using data for a polymer with $N=2048$. The data for different $t$ show a reasonable data collapse. The data in the regime $x\ll 1$ show an increase consistent with the power law $\sim x^{2}$ as shown by the solid line, which indicates that possibly $\tau=0$. Thus one can expect that the validity of the scaling relation Eq. (12). Recalling $z=1.67(3)$ and $w=3.21(25)$, it can be inferred that within numerical error for flexible polymer indeed that is the case. Overall, we conclude that colloid-like dynamic scaling for cluster-cluster aggregation can also be realized in case of clusters composed of monomers during collapse of a flexible polymer.
@@ IV.2 Semi-flexible polymers
Representative evolution snapshots and corresponding contact maps during the collapse of a semiflexible polymer of length $N=2048$ with bending stiffness $\kappa=8$ are presented in Fig. 6. At early times, there is development of clusters, leading to few contacts close to the diagonal. This is followed by aggregation of the small clusters forming of larger regions of off-diagonal contacts. Eventually, a globule is formed where long-range contacts are present over the whole chain. The progressions are similar to those in flexible polymers. However, significantly more order with diamond-like moieties can be noticed in the contact maps for the semi-flexible case. Such order has been reported earlier, and the diamond-like structures in contact maps have been attributed to the ordered local structures for semi-flexible polymers Lappala and Terentjev (2013).
@@
Having an idea about the phenomenology of cluster-cluster aggregation during collapse of semiflexible polymers, we move on to quantify the cluster growth as a function of time. Corresponding plots for different $\kappa$ for a fixed chain length $N=2048$ are presented in Fig. 7(a). As in the case of flexible polymers, $\langle C_{s}(t)\rangle$ for all $\kappa$ display a relatively slow rise at early time, followed by a power-law growth in the scaling regime. As $\kappa$ increases the entry into the scaling regime gets delayed. This can be further confirmed from corresponding plots for time dependence of the number of clusters $n_{c}(t)$ for $\kappa=5$ and $\kappa=8$, presented in the inset of Fig. 7(a). There, the non-monotonic profile for $\kappa=8$ appears because of the time-independent size cut-off, $s_{c}=10$ used to identify the clusters. If one compares the data for the cluster growth for different $\kappa$, it appears that in the scaling regime they are almost parallel to each other, implying the presence of a universal growth exponent $z$, irrespective of $\kappa$. Nevertheless, using the ansatz in Eq. (16) we calculate the average exponent $\langle z\rangle$ within the scaling regime as
$\langle z\rangle=\left\langle\frac{d\ln\left(\langle C_{s}(t)\rangle-C_{0}\right)}{d\ln t}\right\rangle.$ (20)
Here the $\langle\dots\rangle$ indicates averaging over sliding time windows within the scaling regime. The obtained results are compiled in Table 1 along the with their standard deviations and the used range $t_{0},t_{\rm max}]$. We have chosen $t_{0}$ as the point where $\langle n_{c}\rangle$ starts to decay following an initial nucleation stage. Since, the nucleation stage gets prolonged as $\kappa$ increases the chosen $t_{0}$s also have the same trend. This can also be appreciated from the plots of cluster growth in Fig. 7(a), which shows prolonged initial plateau with increasing $\kappa$. The results yield no systematic dependence of the estimated $z$ values on $\kappa$. Rather it fluctuates around a mean value of $z=1.67$. Of course, to further confirm the presence of a $\kappa$-independent $z$, one can rely on finite-size scaling analysis of individual $\kappa$, as done for the flexible case in the previous subsection. However, for the sake of brevity, here, we do not present such exercises for different $\kappa$.
Table 1: Values of mean cluster growth exponent $\langle z\rangle$ obtained by using (20) for polymers with different stiffness $\kappa$.
$\kappa$ $C_{0}$ $\langle z\rangle$ $t_{0},t_{\rm max}]$
0 $32.0$ $1.68(14)$ $[10:140]$
11 $19.0$ $1.68(09)$ $[20:150]$
22 $18.0$ $1.63(06)$ $[25:160]$
33 $17.0$ $1.63(04)$ $[25:200]$
44 $16.0$ $1.67(05)$ $[30:250]$
55 $15.0$ $1.68(06)$ $[40:280]$
66 $8.0$ $1.68(09)$ $[40:300]$
77 $8.0$ $1.68(10)$ $[50:320]$
88 $6.0$ $1.68(13)$ $[50:330]$
99 $3.0$ $1.67(19)$ $[50:350]$
$10$ $3.0$ $1.66(20)$ $[50:370]$
@@
Figure 8: Verification of dynamic scaling in semiflexible polymers. Double-log plot of the scaling function $f(x)\equiv\langle s^{2}N_{s}(t)\rangle$ against the scaling variable $x\equiv s/t^{z}$ at different times for a polymer of length $N=2048$ with bending stiffness (a) $\kappa=5$ and (b) $\kappa=8$. The dashed black lines show the consistency of the data with the power law $\sim x^{\langle\delta\rangle}$ in the regime $x\ll 1$, with respective values of $\langle\delta\rangle$ as indicated. Collapse of data at different times confirm the presence of colloid-like dynamic scaling in semiflexible polymers.
@@
Table 2: Estimates of $\langle w\rangle$ using Eq. (21) to the data for decay of the number of cluster $N_{s}(t)$ of a fixed size $s$ as a function of time for polymers with $N=2048$ and different $\kappa$.
$\kappa$ $\langle w\rangle$ $t_{0},t_{\rm max}]$
0 $3.21(25)$ $[10,140]$
11 $3.05(27)$ $[20,150]$
22 $3.26(26)$ $[25,160]$
33 $3.25(17)$ $[25,200]$
44 $3.18(13)$ $[30,250]$
55 $2.88(28)$ $[40,280]$
66 $2.78(38)$ $[40,300]$
77 $2.73(17)$ $[50,320]$
88 $2.15(11)$ $[50,330]$
99 $2.03(40)$ $[50,350]$
$10$ $1.56(19)$ $[50,370]$
@@
To obtain the exponent $w$, we probe the time dependence of $N_{s}(t)$, i.e., by calculating the variations in the count of clusters of size $s$ with time. Representative plots for $N=2048$ and $s=40$ is presented in Fig. 7(b). As $\kappa$ increases one notices a progressively prominent bell-shaped non-monotonic behavior for the same value of $s$. This is due to the fact that with increasing $\kappa$ nucleation of clusters to form stable clusters gets delayed, as also evident from pots of number of clusters $\langle n_{c}(t)\rangle$ in the inset of Fig. 7(a). For each considered $\kappa$, using the expression in Eq. (19) we calculate the average exponent as
$\langle w\rangle=\left\langle-\frac{d\ln N_{s}(t)}{d\ln t}\right\rangle.$ (21)
Here again the $\langle\dots\rangle$ represents average over sliding time windows. The obtained estimates of $w$ along with their standard deviations are quoted in Table 2 for different $\kappa$. The $t$-range used are chosen to be the same as in Table 1. The value of $w$ appears to decrease with increasing $\kappa$. Also, the value of $w$ becomes smaller than $2z$ as the polymer becomes stiffer. This suggests deviation from the scaling relation in of Eq. (12). The deviations motivate us to ask whether the dynamic scaling at all exists for semi-flexible polymers. If yes, then possibly the cluster-cluster aggregation in semiflexible polymer can be represented by the other kind of dynamic scaling functional form as embedded in Eqs. (8) and (10).
@@
Table 3: Estimates of $\langle\delta\rangle$ using Eq. (23) to the data of the scaling function $f(x)$ as a function of the scaling variable $x$ at different times for polymers with fixed length $N=2048$ and different $\kappa$.
$\kappa$ $\langle\delta\rangle$ $x$-range
0 $2.03(08)$ $[0.001,0.3]$
11 $2.00(20)$ $[0.001,0.2]$
22 $1.92(13)$ $[0.001,0.2]$
33 $2.06(06)$ $[0.001,0.15]$
44 $1.94(24)$ $[0.001,0.12]$
55 $1.91(13)$ $[0.001,0.08]$
66 $1.77(19)$ $[0.001,0.07]$
77 $1.74(12)$ $[0.001,0.06]$
88 $1.64(10)$ $[0.001,0.05]$
99 $1.57(25)$ $[0.001,0.04]$
$10$ $1.48(28)$ $[0.001,0.03]$